OpenStudy (anonymous):

Determine which two values the following number is between. the square root of 176 cubed 5.62 and 5.63 5.6 and 5.61 5.61 and 5.62 5.59 and 5.6

OpenStudy (anonymous):

order of operations is important does the problem read \[(\sqrt{176})^3\] ??

OpenStudy (anonymous):

no parentheses and the 3 is in the front

OpenStudy (anonymous):

do you mean the cubed root?\[\sqrt[3]{176}\] ??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[\sqrt[3]{176}\] = 5.604 The answer is between 5.6 and 5.61 Also, please try to use the equation editor. It will make it easier for others to help you.

OpenStudy (anonymous):

THANKS! <3

OpenStudy (anonymous):

you're welcome if you feel that i have answered your question, you can click best response that would be helpful, thanks